How do you solve -12p=30p? I think you factor but how do you do that when you have only one or two numbers, because if im correct you subtract 30 from both sides to get 12p-30=0, but if you're trying to factor after that how do you do so? Please provide a step by step on this for me. Thank you!
Answer by lynnlo(4176)

-12P=30P
-12P-30P=0
(-12-30)P
(-42)P=0
-42P=0
-42(1)=0
-42.1=0
-42=0====================RESULT
P=0==================SOLUTION